OpenStudy (anonymous):

If ABC is a triangle with AB=20,BC=22 and CA=24. Let D lie on BC such that AD is the angle bisector of ∠BAC.

OpenStudy (anonymous):

find \[AB^2\]

OpenStudy (anonymous):

is an angle bisector also a side bisector

OpenStudy (anonymous):

ie in a triangle

OpenStudy (anonymous):

@satellite73 @shubhamsrg @phi

OpenStudy (phi):

in general, no the angle bisector does not bisect the opposite side (It does if you had an isosceles triangle)

OpenStudy (phi):

what are you trying to find ?

OpenStudy (anonymous):

AB^2

OpenStudy (phi):

Does AB mean length of side AB ? Because they give you that, and you just square it.

OpenStudy (anonymous):

sorry AD^2

OpenStudy (phi):

This looks painful One way is to use the law of cosines and find all three angles A, B, and C of the triangle. Can you do that ?

OpenStudy (phi):

once you have the 3 angles we can take the next step.

OpenStudy (anonymous):

wat i did so far \[x^2=AD^2+24^2-2AB24\cos \theta\] \[(22-x)^2=AD^2+20^2-2AD20\cos \theta\] \[22^2=24^2+20^2-2(24)(20)\cos 2\theta\]

OpenStudy (anonymous):